# Understanding the chemistry of temperate exoplanets atmospheres through experimental and numerical simulations

### v1

@@ 3 ENS-Paris-Saclay, Gif sur Yvette, France
Understanding the chemistry of temperate exoplanets atmospheres through experimental and numerical simulations
@@
(Received ??; accepted ??)
@@
Methods: Our approach combines experimental and numerical simulations on H2-rich gas mixtures representative of sub-Neptune atmospheres, spanning a wide range of CH3, CO, and CO2 mixing ratios. A cold plasma reactor simulates out-of-equilibrium upper-atmospheric chemistry. Chemical evolution is tracked by Mass Spectrometry (MS) and InfraRed (IR) spectroscopy. A 0D photochemical model reproduces reactor conditions, guiding interpretation of key pathways and abundance trends.
Results: We observe the formation of both reduced and oxidized organic compounds. In CH4-rich mixtures, hydrocarbons form efficiently through methane chemistry, correlating with CH4 concentration and agreeing with models. In more oxidizing environments, particularly CO2-rich mixtures, hydrocarbon formation is inhibited by complex reaction networks and oxidative losses. Oxygen incorporation enhances chemical diversity and promotes formation of oxidized organic compounds of prebiotic interest (H2CO, CH3OH, CH3CHO), especially in atmospheres containing both CH4 and CO2. Atmospheres containing CH4 and CO, which balance carbon and oxygen supply without excessive oxidative destruction, favor efficient production of hydrocarbons and oxidized compounds.
@@ 1 Introduction
Characterizing the atmospheres of temperate exoplanets presents significant challenges due to their small size and relatively cold temperatures, making them difficult to detect and study. These exoplanets, which have an equilibrium temperature at below $\sim$500 K, can be classified into three categories: terrestrial planets, super-Earths, and sub-Neptunes. To date, no terrestrial exoplanets, including those in the TRAPPIST-1 system, have been clearly and robustly detected as having an atmosphere (Gillon_2025 for TRAPPIST-1b and c, Piaulet_2025 for TRAPPIST-1d, Espinoza_2025; Glidden_2025 for TRAPPIST-1e). On the other hand, sub-Neptunes and super-Earths — which, to this day, are the most commonly detected types, and therefore dominate the population of known exoplanets (Fressin_2013; Luger_2015; Owen_2016; VanEylen_2018; Hardegree_2019) — have emerged as ideal targets for atmospheric characterization. In particular, those orbiting M dwarf stars have a unique advantage known as the “M dwarf opportunity,” because the smaller size and lower brightness of the host star improve the signal-to-noise ratio during observations. This results in larger transit depths, making it easier to detect and study planetary atmospheres (Cabot_2024; Doyon_2024). These planets are also particularly intriguing because they have no direct analogues in our solar system. This raises fundamental questions about the processes that shape their atmospheres, interiors, formation, and habitability (Cabot_2024). In addition, a significant number of planets orbiting M-dwarfs reside in their star’s temperate zone (Dressing_2015), making the question of their habitability all the more interesting. This habitability is challenged by the unique environmental conditions around M-dwarfs, including intense stellar flares, high-energy particle events, and limited ultraviolet flux to drive prebiotic chemistry (Scalo_2007; Ricker_2015; Shields_2016; Rimmer_2018).
@@
Recent advancements in observational techniques, particularly with instruments on the James Webb Space Telescope (JWST), have made these planets favorable targets. Colder exoplanets between 500 and 1000 K tend to have clearer atmospheres, being less obscured by clouds and hazes (Morley_2015; Brande_2024; Yang_2024; Gao_2021). These advancements have already led to the successful characterization of two temperate sub-Neptune exoplanetary atmospheres, marking significant progress in this field. The analysis of transit spectroscopy data from the atmospheres of K2-18 b (Madhusudhan_2023) and TOI-270 d (Benneke_2024; Holmberg_2024) has highlighted atmospheric compositions dominated by hydrogen, and comprising a few percent of carbon-bearing species such as methane or carbon dioxide. However, despite these revolutionary new data, their atmospheric composition is still poorly constrained and highly debated.
@@
In the case of K2-18 b, initial JWST observations in the 0.9–5.2 µm range, combine data from NIRISS SOSS and NIRSpec G235H instruments acquired in 2023, suggesting the presence of CH4 and CO2 at comparable levels, around 1% each (Madhusudhan_2023). Subsequent studies incorporating non-equilibrium chemistry led to a revised interpretation, indicating a higher level of CH4 and a lower abundance of CO2 (Wogan_2024). This reinterpretation is supported by a new analysis of the JWST data, which has called into question the initial detection of CO2 as a whole (Schmidt_2025). In 2024 and early 2025, four new transit observations were acquired with the NIRSpec G395H instrument and presented by Hu_2025. By stacking all NIRSpec spectral data obtained to date, they report a significant detection of both CH4 and CO2, with upper limits of approximately 10% and 1% respectively, supporting the hypothesis of a predominance of methane. Meanwhile, first data were also acquired with the JWST MIRI LRS instrument, in the 6–12 µm range. Their primary analysis did not provide strong evidence for CH4 or CO2 (Madhusudhan_2025), which do not show clear bands in this range.
@@
The detection of signatures of minor atmospheric constituents also remains largely unconstrained by these low-resolution observations. Non-equilibrium chemistry models help to characterize the complete atmospheric composition, but they require observational constraints such as metallicities and C/O ratios, which are also poorly known. They predict the abundance of various components, even minor ones for which signatures remain undetectable in observations. According to these simulations, CH4, CO2, CO, N2, NH3, H2O, C2H4, C2H6 could be major species, up to the percent range (Jaziri_2025; Luque_2025). Jaziri_2025 also predict that oxidized species such as H2CO, CH3OH and CH3CHO can reach mixing ratios close to the ppm level. These molecules are of significant prebiotic interest. H2CO is important as a possible precursor to sugars through the formose reaction, that is key for the origin of the RNA or pre-RNA world (Butlerow_1861; Cleaves_2008). CH3OH is a key synthon in the RNA world (Mathew_2022). Aldehydes, such as H2CO and CH3CHO, are key precursors of several amino-acids through the Strecker synthesis (Strecker_1850). The atmospheric photochemical production of these oxidized compounds has already been identified as one of the main sources of these compounds in the atmosphere of primitive Earth (Pinto_1980; Junshan_Pinto_1989). It is therefore worthwhile to study their formation pathways in temperate exoplanetary atmospheres.
@@
If we are to overcome the major scientific challenge of characterizing the chemistry of temperate exoplanet atmospheres, a synergy between the observational, modeling and experimental communities is required. Initiatives of this type are already underway in the community. One example is the attempt to obtain experimental data on the temperature-dependent photo-absorption cross sections of key molecules such as CO2, which provide essential data for atmospheric models of hot exoplanets (Venot_2013; Venot_2018). For the cooler exoplanets, studies have been carried out on the detectability of ions in the thermosphere of sub-Neptunes by combining laboratory experiments, chemical modeling and simulated observations, and in the atmospheres of super-Earths by combining laboratory measurements and a 0D photochemical model (Bourgalais_2020; Bourgalais_2021).
@@
A few experiments have been developed on the characterization of the atmospheres of exoplanets, focusing mainly on the formation and properties of haze and aerosols in the atmospheres of super-Earths and sub-Neptunes (Gavilan_2018; He_2018; He_Horst_2018_lab; Horst_2018; Moran_2020; Escobar_2025) or hot Jupiters (Fleury_2019; Fleury_2020; Fleury_2023). Studies focusing on the neutral gas phase chemistry in the atmospheres of cool exoplanets are quite rare. He_Horst_2019_gas conducted laboratory experiments on a wide range of exoplanet atmospheres, with a wide variety of metallicities (100×, 1000×, and 10000× solar metallicity) and temperatures (at 300, 400 and 600 K), and for warmer H2-rich atmospheres (800 K; 100x and 1000x solar metallicities (He_2020)). They explored the effect of different types of irradiation, using a cold plasma or a UV photon lamp as energy sources. They found that a complex non-equilibrium chemistry occurred, leading to the formation of potential haze-forming precursors. Wang_2025 explored CO2-rich atmospheres at 300 and 500 K, with a 2000x solar metallicity. Under plasma irradiation, they observed the formation of various hydrocarbons, oxygen- and nitrogen-containing species, and identified reactive gas precursors for haze formation such as C2H4, H2CO, and HCN. However, the analytical diagnosis of all these studies was based solely on residual gas analyzer mass spectrometry (RGA MS), which led to an uncertain identification of some compounds. Reed_2024 studied the UV irradiation of H2S/CH4/N2 gas mixtures, with or without CO2, at ambient temperature (around 300 K). They analyzed the gas-phase products using gas chromatography equipped with sulfur chemiluminescence detection (GC-SCD), a less ambiguous diagnostic tool than RGA MS. They were thus able to demonstrate the abiotic production of several organosulfur gases, limiting their robustness as biosignatures.
@@
The aim of this work is to propose chemical mechanisms occuring in a wide range of possible atmospheric compositions of temperate sub-Neptunes. In this context, we have carried out parallel experimental and 0D numerical simulations to understand the complex out-of-equilibrium chemistry occurring in the upper layers of the H2-dominated atmospheres of temperate exoplanets. These low-pressure atmospheric layers correspond with those probed by JWST observations and are exposed to energetic radiation that triggers photochemical processes. As the atmospheric composition of these exoplanets is unconstrained, we carried out a study on various H2/CH4/CO/CO2 gas mixtures. The C/O ratio (from 0.5 to infinity) and metallicity (from 5 to 50 times the solar metallicity) of these mixtures are variable. The C/O ratio measures the relative abundance of carbon compared to oxygen, while metallicity measures the abundance of elements heavier than hydrogen and helium. To vary these parameters in our H2-dominated gas mixtures, we use either CH4, CO or CO2 as the main component, as well as different combinations of these carbon compounds. We have chosen to focus on these reactants that allow us to obtain a wide range of possible atmospheres, based on the observational data currently available for K2-18 b. Of course, these are relatively simplistic atmospheric analogues that do not take into account other elements such as nitrogen or sulfur, which may also be present in this type of atmosphere. Other oxygen-bearing molecules, such as water, could also be present in significant quantities in this atmosphere, although the upper layers may instead be dried out by cold trap processes (Charnay_2021). Nevertheless, these atmospheric analogues already allow us to cover a wide range of atmospheric possibilities and to identify a great molecular diversity. A key aspect of this work lies in the combination of experiments and numerical simulations, which enables us to extract abundance trends and chemical formation pathways. Identifying key mechanisms enables us to understand the role of different carbon-bearing reactants (CH4, CO, CO2) in organic growth processes, i.e. carbon chains elongation, and to study the effect of oxygen incorporation on chemical diversification.
@@ 2 Methods
We studied the chemical evolution of atmospheric analogue mixtures when irradiated in a cold plasma setup, mimicking photochemistry in the upper atmosphere. Complementary diagnoses were performed using mass spectrometry (MS) and infrared spectroscopy (IR). In parallel, a 0D photochemical model predicted abundance values and reaction pathways, which we systematically compared with experimental results.
2.1 Mixtures of atmospheric analogues
Different gas mixtures were chosen to study a wide range of potential atmospheres, based on information gathered from observations of K2-18 b. These mixtures are dominated by hydrogen, with carbon-bearing compounds ranging from highly reduced carbon (CH4) to an intermediate state (CO) to fully oxidized carbon (CO2). To enable comparisons with planetary atmospheric models, the C/O ratios and metallicities of our mixtures are provided for reference. It should be noted that our compositions are set by the continuous gas flows in an open system. The gas residence time, i.e. the time spent by the gas in the irradiated medium, is set by the pressure and gas flow in our experiments and thus differ from the residence time found in atmospheric models. Thus, the C/O ratios given are enriched in CH4, CO and CO2, compared to atmospheric models. All mixtures are summarized in Table 1.
The most reduced atmospheres are composed of H2 and CH4. The relative proportions of H2 and CH4 vary in the experiments (see Table 1 for the relative mixing ratios). These mixtures will be referred to as reduced mixtures. This type of atmosphere helps us characterize organic growth in the most reducing environment.
The more oxidized atmospheres are composed of H2 and CO or CO2. The relative mixing ratios are shown in Table 1. These mixtures will be referred to as oxidized mixtures. This type of atmosphere allows us to explore the balance between adding carbon and oxygen atoms to the system.
@@ 2.2 Atmospheric simulation chamber
Experimental simulations of the upper atmospheric layers of temperate exoplanets were carried out using an experimental cold plasma device, the PAMPRE reactor (for Production d’Aérosols en Microgravité par Plasma REactif (Szopa_2006)). This setup was initially developed to simulate the atmospheric chemistry of Titan, and the production of aerosol analogues (tholins) created by photochemistry in this atmosphere (Szopa_2006; Quirico_2008; Sciamma_2010; Carrasco_2012; Mahjoub_2012; Fleury_2014; Dubois_2019; Dubois_2020). The use of this experimental platform has been extended over the years to include other atmospheres, such as early Earth (Fleury_2015; Fleury_2017) or Pluto (Jovanovic_2021). More generally, the versatility of the experimental set-up means we can study any temperate atmospheric chemical composition, which make it suitable for studying exoplanets. Initial research has already been carried out on organic aerosols produced in the atmospheres of Earth-like exoplanets (Gavilan_2017; Gavilan_2018).
@@
PAMPRE induces a cold capacitively coupled radiofrequency plasma in a gaseous mixture, outside of thermodynamic equilibrium. This reproduces the energetic processes occurring in the upper layers of a temperate atmosphere exposed to energetic stellar particles and UV irradiation. It has been previously shown that the energy distribution of electrons in a nitrogen-dominated plasma reproduces the energy distribution of solar photons fairly accurately, with a slightly higher contribution in the VUV (Szopa_2006). The stellar spectra of M dwarfs — around which the temperate exoplanets that interest us orbit — have higher UV radiation contributions than the Sun (France_2013). This experimental setup is therefore appropriate for simulating the energetic environment of these exoplanetary atmospheres. Although our plasma is dominated by hydrogen, which may slightly alter the energy distribution of electrons, it remains adequate. These energetic processes trigger an imbalance in chemistry, dissociating gas molecules and allowing fragments to reorganize into new species. In specific cases, products even aggregate into solid particles, the tholins.
@@
The set-up consists of a cylindrical stainless steel reactor, 40 cm high and 30 cm in diameter, connected to a quadrupole mass spectrometer (QMS). High-pressure gas cylinders are connected to the set-up via an injection manifold, enabling gas mixtures to be introduced into the main chamber. MKS mass flow meters, rated at 5, 10 and 100 standard cubic centimeters per minute (sccm) with a full-scale accuracy of 1%, control the injection rates. The plasma is confined in a stainless steel cage measuring 13.8 cm in diameter and 5 cm in height, pierced with small holes to allow the gas to pass through. The top of this cage is the cathode, measuring 12.6 cm in diameter, polarized and powered by a 13.56 MHz generator. The gas mixture is injected into the reactor from the top of the chamber through the cathode, which is equipped with three superimposed grids to ensure a continuous, homogeneous injection flow. Gas flow rate is set at 40 sccm. The central grid, with a mesh size in the order of the Debye length ($\sim$0.1 mm), prevents the plasma from flowing back into the gas injection system. The gas mixture is extracted from the bottom of the chamber by a primary vane pump (Adixen by Pfeiffer Vacuum Pascal 2015 SD). This enables an open continuous-flow system, at 0.98 mbar and at ambient temperature ($\sim$300 K). A window allows visual monitoring of the plasma. Plasma discharge is only used in continuous mode in these experiments, with a power of 30 W. Between each experiment, the reactor is pumped with a turbomolecular pump (Adixen ATP 80) to achieve a secondary vacuum of around 1.5×10-6 mbar. A schematic diagram of the experimental set-up is shown in Figure 1.

### v2

# Understanding the chemistry of temperate exoplanet atmospheres through experimental and numerical simulations

@@ 3 ENS-Paris-Saclay, Gif sur Yvette, France
Understanding the chemistry of temperate exoplanet atmospheres through experimental and numerical simulations
@@
(Received 03 October 2025; accepted 17 December 2025)
@@
Methods: Our approach combines experimental and numerical simulations on H2-rich gas mixtures representative of sub-Neptune atmospheres and spanning a wide range of CH4, CO, and CO2 mixing ratios. We used a cold plasma reactor to simulate out-of-equilibrium upper-atmospheric chemistry. Chemical evolution was tracked by mass spectrometry and infrared spectroscopy. A 0D photochemical model was used to reproduce reactor conditions, guiding interpretation of the key pathways and abundance trends.
Results: We observed the formation of both reduced and oxidized organic compounds. In CH4-rich mixtures, hydrocarbons formed efficiently through methane chemistry, correlating with CH4 concentration and agreeing with models. In more oxidizing environments, particularly CO2-rich mixtures, hydrocarbon formation was inhibited by complex reaction networks and oxidative losses. We find that oxygen incorporation enhances chemical diversity and promotes the formation of oxidized organic compounds of prebiotic interest (H2CO, CH3OH, CH3CHO), especially in atmospheres containing both CH4 and CO2. Atmospheres containing CH4 and CO — which balance carbon and oxygen supply without excessive oxidative destruction — favor efficient production of hydrocarbons and oxidized compounds.
@@ 1 Introduction
Characterizing the atmospheres of temperate exoplanets presents significant challenges due to their small size and relatively cold temperatures, making them difficult to detect and study. These exoplanets, which have an equilibrium temperature below $\sim$500 K, can be classified into three categories: terrestrial planets, super-Earths, and sub-Neptunes. To date, no terrestrial exoplanets, including those in the TRAPPIST-1 system, have been clearly and robustly detected as having an atmosphere (Gillon et al. (2025) for TRAPPIST-1b and c, Piaulet-Ghorayeb et al. (2025) for TRAPPIST-1d, Espinoza et al. (2025); Glidden et al. (2025) for TRAPPIST-1e). On the other hand, sub-Neptunes and super-Earths — which to this day are the most commonly detected types and therefore dominate the population of known exoplanets (Fressin et al., 2013; Luger and Barnes, 2015; Owen and Mohanty, 2016; Van Eylen et al., 2018; Hardegree-Ullman et al., 2019) — have emerged as ideal targets for atmospheric characterization. In particular, those orbiting M dwarf stars have a unique advantage known as the “M dwarf opportunity,” because the smaller size and lower brightness of the host star improve the signal-to-noise ratio during observations. This results in larger transit depths, making it easier to detect and study planetary atmospheres (Cabot et al., 2024; Doyon, 2024). These planets are also particularly intriguing because they have no direct analogs in our Solar System. This raises fundamental questions about the processes that shape their atmospheres, interiors, formation, and habitability (Cabot et al., 2024). In addition, a significant number of planets orbiting M dwarfs reside in their star’s temperate zone (Dressing and Charbonneau, 2015), making the question of their habitability all the more interesting. This habitability is challenged by the unique environmental conditions around M dwarfs, including intense stellar flares, high-energy particle events, and limited ultraviolet flux to drive prebiotic chemistry (Scalo et al., 2007; Ricker et al., 2015; Shields et al., 2016; Rimmer et al., 2018).
@@
Recent advancements in observational techniques, particularly with instruments on the James Webb Space Telescope (JWST), have made these planets favorable targets. In particular colder exoplanets, with temperatures between 500 and 1000 K, tend to have clearer atmospheres, as they are less obscured by clouds and hazes (Morley et al., 2015; Brande et al., 2024; Yang and Hu, 2024; Gao et al., 2021). The advancements have already led to the successful characterization of two temperate sub-Neptune exoplanetary atmospheres, marking significant progress in this field. The analysis of transit spectroscopy data from the atmospheres of K2-18 b (Madhusudhan et al., 2023) and TOI-270 d (Benneke et al., 2024; Holmberg and Madhusudhan, 2024) has highlighted atmospheric compositions dominated by hydrogen and containing a few percent of carbon-bearing species, such as methane or carbon dioxide. Despite these revolutionary new data, their atmospheric composition is still poorly constrained and highly debated.
@@
In the case of K2-18 b, initial JWST observations in the 0.9–5.2 µm range combined with data from NIRISS SOSS and NIRSpec G235H instruments acquired in 2023 suggest the presence of CH4 and CO2 at comparable levels, around 1% each (Madhusudhan et al., 2023). Subsequent studies incorporating non-equilibrium chemistry have led to a revised interpretation, indicating a higher level of CH4 and a lower abundance of CO2 (Wogan et al., 2024). This reinterpretation is supported by a new analysis of the JWST data, which has called into question the initial detection of CO2 as a whole (Schmidt et al., 2025). In 2024 and early 2025, four new transit observations were acquired with the NIRSpec G395H instrument and presented by Hu et al. (2025). By stacking all NIRSpec spectral data obtained to date, they report a significant detection of both CH4 and CO2, with upper limits of approximately 10% and 1% respectively, supporting the hypothesis of a predominance of methane. Meanwhile, initial data were also acquired with the JWST MIRI LRS instrument in the 6–12 µm range, and primary analysis of the data did not provide strong evidence for CH4 or CO2 (Madhusudhan et al., 2025), which do not show clear bands in this range.
@@
The detection of signatures of minor atmospheric constituents also remains largely unconstrained by these low-resolution observations. Non-equilibrium chemistry models help characterize the complete atmospheric composition, but they require observational constraints such as metallicities and C/O ratios, which are also poorly known. They predict the abundance of various components, even minor ones for which signatures remain undetectable in observations. According to these simulations, CH4, CO2, CO, N2, NH3, H2O, C2H4, and C2H6 could be major species, up to the percent range (Jaziri et al., 2025; Luque et al., 2025). Jaziri et al. (2025) also predict that oxidized species such as H2CO, CH3OH, and CH3CHO can reach mixing ratios close to the parts-per-million level. These molecules are of significant prebiotic interest. H2CO is important as a possible precursor to sugars through the formose reaction, which is key for the origin of the RNA or pre-RNA world (Butlerow, 1861; Cleaves, 2008). CH3OH is a key synthon in the RNA world (Mathew et al., 2022). Aldehydes, such as H2CO and CH3CHO, are key precursors of several amino acids through the Strecker synthesis (Strecker, 1850). The atmospheric photochemical production of these oxidized compounds has already been identified as one of the main sources of these compounds in the atmosphere of primitive Earth (Pinto et al., 1980; Wen et al., 1989). It is therefore worthwhile to study their formation pathways in temperate exoplanetary atmospheres.
@@
If we are to overcome the major scientific challenge of characterizing the chemistry of temperate exoplanet atmospheres, a synergy between the observational, modeling, and experimental communities is required. Initiatives of this type are already underway in the community. One example is the attempt to obtain experimental data on the temperature-dependent photo-absorption cross sections of key molecules such as CO2, which provide essential data for atmospheric models of hot exoplanets (Venot et al., 2013, 2018). For the cooler exoplanets, studies have been carried out on the detectability of ions in the thermosphere of sub-Neptunes by combining laboratory experiments, chemical modeling, and simulated observations as well as in the atmospheres of super-Earths by combining laboratory measurements and a 0D photochemical model (Bourgalais et al., 2020, 2021).
@@
A few experiments have been developed for the characterization of the atmospheres of exoplanets, and they mainly focus on the formation and properties of haze and aerosols in the atmospheres of super-Earths and sub-Neptunes (Gavilan et al., 2018; He et al., 2018a, b; Hörst et al., 2018; Moran et al., 2020; Jiménez-Escobar et al., 2025) or hot Jupiters (Fleury et al., 2019, 2020, 2023). Studies focusing on the neutral gas phase chemistry in the atmospheres of cool exoplanets are quite rare. He et al. (2019) conducted laboratory experiments on a wide range of exoplanet atmospheres, with a wide variety of metallicities (100×, 1000×, and 10000× solar metallicity) and temperatures (at 300, 400, and 600 K) and for warmer H2-rich atmospheres (800 K, 100x, and 1000x solar metallicities (He et al., 2020)). They explored the effect of different types of irradiation, using a cold plasma or a UV photon lamp as energy sources. They found that a complex non-equilibrium chemistry occurred, leading to the formation of potential haze-forming precursors. Wang et al. (2025) explored CO2-rich atmospheres at 300 and 500 K, with a 2000x solar metallicity. Under plasma irradiation, they observed the formation of various hydrocarbons and oxygen- and nitrogen-containing species, and they identified reactive gas precursors for haze formation such as C2H4, H2CO, and HCN. However, the analytical diagnosis of all these studies was based solely on residual gas analyzer mass spectrometry (RGA MS), which led to an uncertain identification of some compounds. Reed et al. (2024) studied the UV irradiation of H2S/CH4/N2 gas mixtures, with or without CO2, at an ambient temperature (around 300 K). They analyzed the gas-phase products using gas chromatography equipped with sulfur chemiluminescence detection (GC-SCD), a diagnostic tool less ambiguous than RGA MS. They were thus able to demonstrate the abiotic production of several organosulfur gases, limiting their robustness as biosignatures.
@@
The aim of this work is to propose chemical mechanisms occurring in a wide range of possible atmospheric compositions of temperate sub-Neptunes. In this context, we carried out parallel experimental and 0D numerical simulations to understand the complex out-of-equilibrium chemistry occurring in the upper layers of the H2-dominated atmospheres of temperate exoplanets. These low-pressure atmospheric layers correspond to those probed by JWST observations and were exposed to energetic radiation that triggers photochemical processes. As the atmospheric composition of these exoplanets is unconstrained, we carried out a study on various H2/CH4/CO/CO2 gas mixtures. The C/O ratio (from 0.5 to infinity) and metallicity (from 5 to 50 times the solar metallicity) of these mixtures are variable. The C/O ratio measures the relative abundance of carbon compared to oxygen, while metallicity measures the abundance of elements heavier than hydrogen and helium. To vary these parameters in our H2-dominated gas mixtures, we used CH4, CO, or CO2 as the main component as well as different combinations of these carbon compounds. We chose to focus on these reactants because they allow us to obtain a wide range of possible atmospheres based on the observational data currently available for K2-18 b. Of course, the reactants are relatively simplistic atmospheric analogs that do not take into account other elements such as nitrogen or sulfur, which may also be present in this type of atmosphere. Other oxygen-bearing molecules, such as water, could also be present in significant quantities in this atmosphere, although the upper layers may instead be dried out by cold trap processes (Charnay et al., 2021). Nevertheless, these atmospheric analogs already allow us to cover a wide range of atmospheric possibilities and to identify a great molecular diversity. A key aspect of this work lies in the combination of experiments and numerical simulations, which enables us to extract abundance trends and chemical formation pathways. Identifying key mechanisms enables us to understand the role of different carbon-bearing reactants (CH4, CO, CO2) in organic growth processes (i.e., carbon chains elongation) and to study the effect of oxygen incorporation on chemical diversification.
@@ 2 Methods
We studied the chemical evolution of atmospheric analog mixtures when irradiated in a cold plasma setup, mimicking photochemistry in the upper atmosphere. Complementary diagnoses were performed using mass spectrometry (MS) and infrared spectroscopy. In parallel, a 0D photochemical model predicted abundance values and reaction pathways, which we systematically compared with experimental results.
2.1 Mixtures of atmospheric analogs
Different gas mixtures were chosen to study a wide range of potential atmospheres, based on information gathered from observations of K2-18 b. These mixtures are dominated by hydrogen, with carbon-bearing compounds ranging from highly reduced carbon (CH4) to an intermediate state (CO) to fully oxidized carbon (CO2). To enable comparisons with planetary atmospheric models, the C/O ratios and metallicities of our mixtures are provided for reference. It should be noted that our compositions are set by the continuous gas flows in an open system. The gas residence time, i.e., the time spent by the gas in the irradiated medium, is set by the pressure and gas flow in our experiments and thus differ from the residence time found in atmospheric models. Thus, the C/O ratios given are enriched in CH4, CO and CO2, compared to atmospheric models. All mixtures are summarized in Table 1.
The most reduced atmospheres are composed of H2 and CH4. The relative proportions of H2 and CH4 vary in the experiments (see Table 1 for the relative mixing ratios). We refer to these mixtures as reduced mixtures. This type of atmosphere helps us characterize organic growth in the most reducing environment.
The more oxidized atmospheres are composed of H2 and CO or CO2. The relative mixing ratios are shown in Table 1. We refer to these mixtures as oxidized mixtures. This type of atmosphere allows us to explore the balance between adding carbon and oxygen atoms to the system.
@@ 2.2 Atmospheric simulation chamber
Experimental simulations of the upper atmospheric layers of temperate exoplanets were carried out using an experimental cold plasma device, the PAMPRE reactor (for Production d’Aérosols en Microgravité par Plasma REactif (Szopa et al., 2006)). This setup was initially developed to simulate the atmospheric chemistry of Titan, and the production of aerosol analogs (tholins) created by photochemistry in this atmosphere (Szopa et al., 2006; Quirico et al., 2008; Sciamma-O’Brien et al., 2010; Carrasco et al., 2012; Mahjoub et al., 2012; Fleury et al., 2014; Dubois et al., 2019, 2020). The use of this experimental platform has been extended over the years to include other atmospheres, such as early Earth (Fleury, 2015; Fleury et al., 2017) or Pluto (Jovanovic, 2021). More generally, the versatility of the experimental setup means we can study any temperate atmospheric chemical composition, which make it suitable for studying exoplanets. Initial research has already been carried out on organic aerosols produced in the atmospheres of Earth-like exoplanets (Gavilan et al., 2017, 2018).
@@
PAMPRE induces a cold capacitively coupled radiofrequency plasma in a gaseous mixture, outside of thermodynamic equilibrium. This reproduces the energetic processes occurring in the upper layers of a temperate atmosphere exposed to energetic stellar particles and UV irradiation. It has been previously shown that the energy distribution of electrons in a nitrogen-dominated plasma reproduces the energy distribution of solar photons fairly accurately, with a slightly higher contribution in the VUV (Szopa et al., 2006). The stellar spectra of M dwarfs — around which the temperate exoplanets that interest us orbit — have higher UV radiation contributions than the Sun (France et al., 2013). This experimental setup is therefore appropriate for simulating the energetic environment of these exoplanetary atmospheres. Although our plasma is dominated by hydrogen, which may slightly alter the energy distribution of electrons, it remains adequate. These energetic processes trigger an imbalance in chemistry, dissociating gas molecules and allowing fragments to reorganize into new species. In specific cases, products even aggregate into solid particles, the tholins.
@@
The setup consists of a cylindrical stainless steel reactor, 40 cm high and 30 cm in diameter, connected to a quadrupole mass spectrometer (QMS). High-pressure gas cylinders are connected to the setup via an injection manifold, enabling gas mixtures to be introduced into the main chamber. MKS mass flow meters, rated at 5, 10 and 100 standard cubic centimeters per minute (sccm) with a full-scale accuracy of 1%, control the injection rates. The plasma is confined in a stainless steel cage measuring 13.8 cm in diameter and 5 cm in height, pierced with small holes to allow the gas to pass through. The top of this cage is the cathode, measuring 12.6 cm in diameter, polarized and powered by a 13.56 MHz generator. The gas mixture is injected into the reactor from the top of the chamber through the cathode, which is equipped with three superimposed grids to ensure a continuous, homogeneous injection flow. Gas flow rate is set at 40 sccm. The central grid, with a mesh size in the order of the Debye length ($\sim$0.1 mm), prevents the plasma from flowing back into the gas injection system. The gas mixture is extracted from the bottom of the chamber by a primary vane pump (Adixen by Pfeiffer Vacuum Pascal 2015 SD). This enables an open continuous-flow system, at 0.98 mbar and at ambient temperature ($\sim$300 K). A window allows visual monitoring of the plasma. Plasma discharge is only used in continuous mode in these experiments, with a power of 30 W. Between each experiment, the reactor is pumped with a turbomolecular pump (Adixen ATP 80) to achieve a secondary vacuum of around 1.5×10-6 mbar. A schematic diagram of the experimental setup is shown in Figure 1.